Calculate the value of K_p for the equation C(s)
Answer #1

Step: 1: write the first equation as it is C + 2H2O--> CO2 + 2H2 Kp1 = 3.75 Step2: multiply the equation by 2 and we will get
Answer #2

This is wrong. You multiply the K values when you add the two equations. 

Kpfinal = Kp1(Kp2)2

When you multiply a reaction by a factor, you have to raise your Kp value to that factor. So above, Kp2= 0.7012=0.4914

When you sum the two reaction equations, you have to multiply your two final Kp's. 

Kp = (0.4914)(3.75) = 1.845